### v1

# Improving MLLMs in Embodied Exploration and Question Answering with Human-Inspired Memory Modeling

Abstract
Deploying Multimodal Large Language Models as the brain of embodied agents remains challenging, particularly under long-horizon observations and limited context budgets. Existing memory-assisted methods often rely on textual summaries, which discard rich visual and spatial details and remain brittle in non-stationary environments. In this work, we propose a non-parametric memory framework that explicitly disentangles episodic and semantic memory for embodied exploration and question answering. Our retrieval-first, reasoning-assisted paradigm recalls episodic experiences via semantic similarity and verifies them through visual reasoning, enabling robust reuse of past observations without rigid geometric alignment. In parallel, we introduce a program-style rule extraction mechanism that converts experiences into structured, reusable semantic memory, facilitating cross-environment generalization. Extensive experiments demonstrate state-of-the-art performance on embodied question answering and exploration benchmarks, yielding a 7.3% gain in LLM-Match and an 11.4% gain in LLM-MatchSPL on A-EQA, as well as +7.7% success rate and +6.8% SPL on GOAT-Bench. Analyses reveal that our episodic memory primarily improves exploration efficiency, while semantic memory strengthens complex reasoning of embodied agents.
I Introduction
In recent years, Multimodal Large Language Models (MLLMs) have advanced rapidly, demonstrating strong reasoning and perception capabilities across a wide range of domains. Despite this progress, deploying MLLMs as the core decision-making module of embodied agents in real-world environments remains highly challenging. A key difficulty lies in enabling agents to efficiently process long-horizon observations while retaining task-relevant information under limited context budgets. To address this issue, recent works introduce various forms working memory, allowing agents to retrieve a small subset of relevant past observations instead of conditioning on the entire sensory history. By coupling task-relevant information retrieval with frontier-based exploration strategies, these methods achieve efficient and effective performance in embodied exploration and embodied question answering tasks.
However, most existing approaches remain confined to information within the current episode, limiting their ability to leverage accumulated experience over time. In contrast, real-world embodied agents are expected to continuously integrate observations across episodes and exploit long-term knowledge acquired from past interactions. While [6] incorporates cross-episode memory, its task definition remain restrictive, typically assessing an MLLM-based information-collection system within a fixed and fully explored space derived from previous timeline. Moreover, approaches such as [30] and [31] fundamentally summarize cross-episode information into hierarchical textual descriptions, which are then used to guide future exploration. This design inevitably discards a substantial amount of spatial information and rich visual cues in raw RGB observations, constraining the agent’s ability to remember.
To eflects real-world embodied scenario, we formulate the question in assumption that environments are non-stationary, and then reveals two fundamental challenges. First, maintaining long-term maps—whether represented as persistent nodes [25] or voxel grids [22]-becomes inherently difficult in the presence of environmental changes, sensor noise, and long-horizon drift. Second, information retrieved from past episodes does not necessarily reflect the current state of the environment. Based on these observations, we argue that an effective embodied agent should neither rely on rigid long-term geometric fusion nor blindly trust past episodic information. Instead, it should recall prior experiences in a soft and associative manner, while continuously grounding them in active exploration and perception — closely mirroring how humans reason about and interact with dynamic environments.
Motivated by this insight, we draw inspiration from human memory systems and propose a non-parametric memory framework that explicitly disentangles episodic and semantic memory for embodied exploration and question answering. First, our approach adopts retrieval-first and reasoning-assisted paradigm, where episodic memories are recalled based on semantic similarity and selectively verified through a system-2 visual reasoning. Second, rather than storing memory as free-form textual summaries which are often verbose, environment-specific, and difficult to reuse, we propose program-style enhanced rule extraction to form semantic memory. By generating pseudocode-style workflows and incorporating decision deviation detection, our approach extracts structured, high-level, and reusable decision rules from embodied experiences. These rules formed semantic memory generalize more effectively across environments, providing stronger and more reliable reasoning support during decision making.
We achieve new state-of-the-art results on embodied question answering and exploration, with 65.6% LLM-Match and 48.7% LLM-MatchSPL on A-EQA [15], and a 72.8% success rate with 56.1% SPL on GOAT-Bench [10]. Extensive experiments show that our proposed episodic memory improves exploration efficiency, while semantic memory enhances performance on complex, long-horizon reasoning tasks.
II Related Works
II-A MLLM-based Embodied Agents
Recent proprietary foundation models, such as GPT-5 and Gemini 3, have significantly advanced multi-image understanding by leveraging extended context lengths and large-scale pretraining [12], and enable more comprehensive perception over long visual sequences, forming a strong backbone for embodied agents [3, 34, 28, 35, 20, 32]. Beyond vision-language foundation models, imagination driven designs augment observations via generative priors [26, 24], while confidence calibration approaches explicitly model when MLLM should stop by quantifying epistemic uncertainty [4, 19, 9].
Despite these advances, foundation model still incur substantial computational costs, and naively processing uniformly sampled frames is often suboptimal for questions that require selectively identifying high-value observations for exploration and answering. To address these challenges, [7, 23, 25, 33] introduces CLIP and SAM-based open-vocabulary scene graphs to structure semantics for MLLM-driven exploration. While [27] further incorporates raw images to enrich memory, its key-frame selection remains constrained by detectors and lacks explicit geometric modeling, leading to inefficient long-horizon exploration. In contrast to existing geometric-semantic graph methods often depend on heavy multi-frame fusion, our approach marks a shift toward decoupling semantic and geometric spaces, enabling retrieval-time interaction that supports zero-shot generalization and adaptive cross-episode recalling.
II-B Experience and Memory for Embodied Agents
Research on memory mechanisms for embodied agents seeks to enable the storage, retention, and retrieval of past experiences, facilitating the transition from purely reactive systems to agents capable of maintaining context and autonomous adaptation. [1] and [30] adopt description text, temporal and spatial cues to manage continuously growing histories, allowing robots to recall previously visited regions and reduce redundant exploration. However, memory built primarily on video or image captioning discard a substantial amount of fine-grained visual information embedded in raw observations. [6] further explores long-term memory under dynamic environments by preserving raw image observations. Nevertheless, its task formulation remains restrictive, typically evaluating an information searching system within a fixed explored space, hinders its applicability to realistic exploration scenarios. More recently, [31] focuses on extracting reusable experience from training trajectories. While effective at capturing high-level patterns, the learned experiences remain tightly coupled to previously environment, limiting their generalization ability. Motivated by these observations, we draw inspiration from human memory systems and propose to disentangle episodic and semantic memory, recalling episode to improve exploration efficiency, while following high-level decision rules in decision making.
III Methods
III-A Semantic and Physical Space for Embodied Agents
Navigation in embodied environments requires reasoning jointly over a semantic space and a physical space. Inspired by cognitive science findings suggesting structured interactions between these two spaces [21], we propose to decouple semantic memory from physical geometry, while enabling their interaction through test time retrieval.
Semantic and Physical Space. At each observation step, we extract a set of region-level semantic embeddings and a place-level image embedding using SAM [11] and CLIP [17] to form hierarchical embedding-based scene graph, following HOV-SG [23]. Global and regional embeddings are stored in a semantic memory indexed by observation time, camera pose and raw image. To interface with downstream tasks, we leverage a LLM to decompose an instruction into three semantic components: a target object, a set of relative objects, and a set of relative areas. Each component is embedded and used to retrieve the most relevant memory via similarity search over region embeddings. Retrieved memory are selected in a priority order, with target object evidence taking precedence, followed by relative objects and relative areas. While the physical space is represented by a lightweight 2D occupancy map constructed from depth observations and camera poses.
Retrieval-Time Inter-Space Interaction. Unlike prior approaches such as HOV-SG, which tightly couple semantic representations with multi-view geometric fusion, our semantic memory is updated independently of persistent geometry. Specifically, we only link semantic space to the physical space at test time by projecting the camera poses of retrieved observations onto the occupancy grid. The interaction allowing semantic concepts to be flexibly grounded without continuous online fusion, and preserves task-relevant semantic cues while keeping the physical representation compact and efficient.
III-B Human-Inspired Memory for Embodied Exploration
Embodied agents are expected to learn and adapt from their past experiences. Based on this, we propose a non-parametric method to model human-inspired memory for embodied exploration and question answering.
| Method | Object Rec. | Object Loc. | Attribute Rec. | Spatial | Object State | Functional | World Know. | Overall | ||||||||
| Socratic LLM-based Exploration w/ Frame Captions | ||||||||||||||||
| GPT-4* | 25.3 | – | 28.4 | – | 27.3 | – | 37.7 | – | 47.2 | – | 54.2 | – | 29.5 | – | 35.5 | |
| GPT-4o | 22.0 | – | 25.0 | – | 27.3 | – | 40.8 | – | 50.9 | – | 61.8 | – | 38.4 | – | 35.9 | |
| Socratic LLM-based Exploration w/ Scene-Graph Captions | ||||||||||||||||
| CG Scene-Graph* | 25.3 | – | 16.5 | – | 29.2 | – | 37.0 | – | 52.2 | – | 46.8 | – | 37.8 | – | 34.4 | |
| SVM Scene-Graph* | 29.0 | – | 17.2 | – | 31.5 | – | 31.5 | – | 54.2 | – | 39.8 | – | 38.9 | – | 34.2 | |
| LLaVA-1.5* | 25.0 | – | 24.0 | – | 34.1 | – | 34.4 | – | 56.9 | – | 53.5 | – | 40.6 | – | 38.1 | |
| Multi-Frame* | 34.0 | – | 34.3 | – | 51.5 | – | 39.5 | – | 51.9 | – | 45.6 | – | 36.6 | – | 41.8 | |
| Open-Sourced MLLM-based Exploration | ||||||||||||||||
[…]
| Commercial MLLM Exploration | ||||||||||||||||
[…]
| Human Agent* | 89.7 | – | 72.8 | – | 85.4 | – | 84.8 | – | 97.8 | – | 78.9 | – | 88.5 | – | 85.1 | |
Episodic Memory: Visual Reasoning for Episodic Recall. Episodic memory records episode-level experiences in a detailed and temporally grounded manner. Each episodic memory instance consists of both a semantic space and a physical space, which encode the observed scene from complementary perspectives, as described in Section III-A. As a fine-grained memory type, episodic memory preserves information at the granularity of navigation steps.
When encountering a new episode, we first retrieve the Top- most similar past observations based on visual similarity to the current observation. To verify episodic correspondence, we further assess whether the retrieved observations originate from nearby locations using an MLLM. After that, embeddings of the current target are used to query the verified episode, and the episode that accumulates the highest number of Top- matches is selected as prior experience.
Mapping-based approaches [13] explicitly align and merge memories by estimating relative transformations, which are highly sensitive to optimization failures and are prone to cascading errors during long-horizon changing and sensor noises. While we humans recall past experiences through associative and semantic similarity rather than rigid metric fusion of memories. Hence, in contrast to mapping methods, we propose a gentle episodic recall strategy supported by visual reasoning capability.
Specifically, given the retrieved Top- episodic observations, MLLM first determines whether further exploration is required for the current task.
Let composes the cognitive map, where represent the retrieved camera poses, and denotes the set of frontier points. For each frontier , let indicate whether it can lead to unexplored regions, and denote its spatial distance to semantically relevant landmarks. The current and previous exploration maps are jointly provided to a MLLM, outputing pruned frontiers in a visual reasoning manner:
Here, another retrieval-time interaction between physical and semantic space is expressed through frontiers pruning operation. Through this reasoning-centric approach, episodic memory informs exploration decisions without requiring explicit geometric fusion, resulting in a stable exploration strategy.
Semantic Memory: Program-Style Enhanced Rule Extraction. Semantic memory captures long-term, consolidated knowledge distilled from an agent’s past experiences, including reusable knowledge and preferred response patterns. Unlike episodic memories, which rely on fine-grained details and are inherently tied to specific environments, our semantic memory aims to extract reusable rules from dilemmas encountered during reasoning execution, with the core objective of capturing solutions that are transferable across environments. However, pioneer experiments reveal that, directly summarizing embodied agents reasoning logs are often overly entangled with environments-specific details that fail to reuse and compactly encode essential knowledge.
To address these limitations, we follow [8] and introduce a structured abstraction strategy based on pseudocode-style workflows. Specifically, we prompt an LLM to analyze the execution trace of a task by prioritizing high-level structural information. The model first identifies variables and functions that govern the workflows, and then organizes them into a pseudocode that explicitly captures control flow and decision logic.
Unlike prior work [8, 14], where workflows are typically short, embodied agent often involves long-horizon execution with extended decision sequences. In this situation, humble LLM assistant tends to over-annotate functions with learned rules that are weakly correlated with task success, obscuring the truly critical decision points. To prevent this, we introduce decision deviation detection mechanism on robots’ trajectories, following [29]. Specifically, given the robot state at time step and the corresponding ground-truth trajectory , we compute:
where measures the Euclidean distance between the agent’s current position and the ground-truth trajectory. Interactively setting a predefined threshold , we identify a decision deviation at timestep if
We randomly stop the detection at threshold when the number of decision deviation is in the range of 3-5 to form set . These deviation timesteps signal potential errors or departures from the expected behavior, allowing us to ‘debug’ the ‘program’.
During the rule extraction, we provide the MLLM with the ground truth answer , , , together with their corresponding visual observations :
Finally, we can abstracts experiential knowledge into structured, high-level rules using canonical forms such as if–then, situation–suggestion and problem–solution key-value pairs, where each pair belongs to one specific variables or functions.
At the test-time, the task-similarity retrieved semantic memory is loaded into the system prompt. Details can be found in the appendix.
Meta Memory: Inspired by the success of cognitive process modeling [2], we implement meta memory via cognitive states automatic switching. Specifically, we pre-define a set of discrete cognitive states that govern high-level reasoning behavior, including: exploration, target verification, double checking, ready to answer. See Table in the appendix for concrete prompt examples associated with each cognitive state.
IV Experiments
IV-A Active Embodied Question Answering
To evaluate the effectiveness of our memory design, we conduct experiments on the Open-EQA benchmark [15, 16, 18]. Here, we compare our approach against other MLLM-based baselines on the Open-EQA benchmark, assessing both exploration efficiency and performance on answering complex open-ended questions. In addition, our semantic memory is distilled from 184 episodes in the training split, with a category distribution aligned with that of the test set.
Metrics. We report LLM-Match and LLM-MatchSPL as evaluation metrics to measure answer quality and exploration efficiency. Specifically, SPL is computed as the ratio between the shortest path length and the actual trajectory length, normalized to , where a higher value indicates a more efficient navigation trajectory. LLM-Match evaluates the semantic correctness of the agent’s answers using GPT-4o as an automated judge. LLM-MatchSPL further weights the semantic correctness score by navigation efficiency, jointly reflecting both answer quality and exploration effectiveness. If the agent fails to produce a response, answer is outputted by GPT-4o without visual input, and the SPL term is set to zero.
Analysis. As shown in Table I, our method consistently outperforms all baseline approaches when using the GPT-4o backbone. Specifically, the overall LLM-Match improves from 58.3% to 65.6%, while LLM-Match×SPL increases substantially from 37.3% to 48.7%, indicating both higher answer accuracy and improved exploration efficiency.
Notably, significant gains are observed across most task categories: Object recognition improves from 45.0% to 62.0%, functional reasoning from 65.0% to 66.4%, and world knowledge from 35.0% to 54.5%. On the one hand, our method achieves higher LLM-Match×SPL across all categories, demonstrating that episodic memory recall guided by visual reasoning effectively directs exploration toward more goal-relevant regions, resulting in more purposeful and efficient navigation. On the other hand, the pronounced improvements in world knowledge and functional recall highlight our semantic memory supporting the reasoning-intensive and cognitively complex tasks. Additionally, strong improvements in object recognition and object localization can be attributed to our framework’s reduced reliance on specialized perception models [5], enabling better and more flexible object identification through semantic space rather than purely textual cues.
When switching to the Qwen-2.5-VL-7B-Instruct backbone, our method maintains consistent performance gains. It achieves the second-best LLM-Match and the best LLM-Match×SPL, indicating a favorable balance between answer accuracy and exploration efficiency across different MLLM backbones. In contrast to ReExplore [31], which employs a stronger MLLM backbone but exhibits noticeable performance degradation in object recognition and world knowledge, our framework continues to scale positively with increasing MLLM capability. This suggests that our design more effectively leverages enhanced MLLM reasoning capacity, rather than being bottlenecked by other components in system.
IV-B Life-Long Visual Language Navigation
GOAT-Bench [10] is a lifelong visual navigation benchmark designed around a multi-target Go-to-AnyThing (GOAT) task. In this setting, an agent is required to navigate sequentially to multiple target objects within previously unseen environments. Targets are specified in diverse modalities, including object category names (e.g., refrigerator), natural language descriptions (e.g., toy next to the dresser), or goal images.
Following prior baseline methods [31, 27], we evaluate our approach on a -scale subset of the val-unseen split, which contains one episode per scene across 36 scenes, resulting in a total of 278 navigation subtasks. In addition, our semantic memory is constructed from the 36 training scenes.
| Method | Success Rate (%) | SPL (%) |
| Traditional Methods | ||
| Modular GOAT [10]* | 24.9 | 17.2 |
| Modular CLIP on Wheels [10]* | 16.1 | 10.4 |
| SenseAct-NN Skill Chain [10]* | 29.5 | 11.3 |
| SenseAct-NN Monolithic [10]* | 12.3 | 6.8 |
| MLLM-based Exploration | ||
| 3D-Mem [27]† (Qwen2.5-VL-7B-Instruct) | 49.4 | 20.7 |
[…]
| Explore-EQA [19]† (GPT-4o) | 55.0 | 37.9 |
| CG w/ Frontier Snapshots [27]† (GPT-4o) | 61.5 | 45.3 |
| 3D-Mem w/o Memory [27]† (GPT-4o) | 58.6 | 38.5 |
[…]
| ReEXplore [31]† (GPT-4o) | 59.8 | 42.5 |
[…]
| Method | Episodic Memory | Semantic Memory | LLM-Match | LLM-MatchSPL |
| Ours (Qwen2.5-VL-7B-Instruct) | ✓ | ✓ | 43.1 | 23.7 |
| ✓ | 41.6 | 23.2 | ||
| ✓ | 39.1 | 17.5 | ||
| Ours (GPT-4o) | ✓ | ✓ | 65.6 | 48.7 |
| ✓ | 62.3 | 47.2 | ||
| ✓ | 60.9 | 42.1 |
Metrics. We adopt Success Rate and Success weighted by Path Length (SPL) as evaluation metrics. An episode is deemed successful when the agent terminates within 1 m of the navigation target. SPL further measures efficiency by weighting successful episodes according to the length of the navigation path.
Analysis. As shown in Table II, our method outperforms most baseline methods, demonstrating its ability to effectively leverage knowledge accumulated from both episode memory and semantic memory, as evidenced by consistently higher Success Rate and SPL scores. In particular, when compared with 3D-Mem under the GPT-4o backbone, our method achieves a notable improvement in life-long navigation, with the Success Rate increasing from 65.1% to 72.8%, and SPL improving from 49.3% to 56.1%, indicating more accurate and efficient goal-directed exploration. Under the Qwen backbone, our method slightly underperforms ReEXplore in SPL metric. We attribute this primarily to the relatively weaker visual reasoning capability of Qwen backbone when operating on semantic-labeled exploration maps, which limits the effectiveness of frontier pruning operation during exploration.
IV-C Ablation Study
Table III presents ablation results on the A-EQA benchmark. When episodic memory is removed, we observe a clear degradation in exploration efficiency. Specifically, LLM-Match drops from 41.3% to 39.1% under the Qwen backbone, and from 65.6% to 60.9% under the GPT-4o backbone. This performance decrease indicates that our episodic memory provides effective guidance for exploration by enabling the agent to reuse relevant prior experiences. In contrast, removing semantic memory results in a more decline in answer accuracy than corresponding reduction in SPL. This highlights the critical role of semantic memory in preserving transferable and high-level rules and supporting embodied reasoning during question answering. Overall, these ablation results demonstrate that the two memory modules offer complementary benefits: episodic memory primarily improves exploration efficiency, while semantic memory contributes more substantially to accurate and reasoning-intensive question answering.
V Conclusion
We presented a non-parametric memory framework for MLLM-based embodied agents that explicitly disentangles episodic and semantic memory to address long-horizon exploration and reasoning. By adopting a visual reasoning paradigm, our approach enables agents to recall and verify prior episodic experiences without relying on rigid geometric fusion, thereby supporting robust cross-episode reuse of observations and improving exploration efficiency. In addition, resuable and structed extracted rules from experiences tackle complex decision making in embodied tasks. Comprehensive evaluation on A-EQA and GOAT-Bench demonstrates the effectiveness of our memory modeling in improving both exploration efficiency and answering accuracy. We identify our limitations points to promising future directions: 1) Differences in scale, map and visual may introduce new challenges for episodic memory in outdoor environments. 2) Parametrically integrating semantic memory into a embodied foundation model for joint optimization of memory recalling.